OpenStudy (anonymous):

I can Help anyone in Algebra 1 B part 1 that goes to Connections academy! :D

OpenStudy (tyrinzoebrown):

oh my god thatnk you I need it. but first I have to solve linear equations on algebra 1 a, do you know how to do that

OpenStudy (anonymous):

Yes but I'll be back in 30 got to go to the store Be back soon! :)

OpenStudy (tyrinzoebrown):

ok ill write my questions just answer when you get the chance

OpenStudy (anonymous):

okay! :D

OpenStudy (tyrinzoebrown):

whats the value for x 1. x-7>10 2. 3x (<or = to)21 3. 5x-2<18 4. 2x+8(>or = to) -4(x+1) 9. sara has to read a book in english class by the end of the week. she wants to read atleast 200 pages in the first 5 days of the week. on adverage how many pages should sara read per day to meet her goal? p (> or = to) ________ 10. A painter earns 15$ an hour. what is the minimum number of hours he must work to earn atleast $200? write an inequality to represent the situation and solve. Show all work!